OpenStudy (kidthatbro8):

Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8^x and y = 2^x + 2 intersect are the solutions of the equation 8^x = 2^x + 2. (4 points) Part B Make tables to find the solution to 8^x = 2^x + 2. Take the integer values of x between -3 and 3. (4 points) Part C: How can you solve the equation 8^x = 2^x + 2 graphically? (2 points)

OpenStudy (igreen):

For Part A..the reason why is because you substituted 8^x for 'y' in the 2nd equation..

OpenStudy (kidthatbro8):

oh, okay.

OpenStudy (anonymous):

21

OpenStudy (igreen):

I'm not sure how we would solve Part B..

OpenStudy (igreen):

Oh, it wants tables from y = 8^x and y = 2^x + 2..

OpenStudy (kidthatbro8):

i don't understand that part...

OpenStudy (igreen):

Here's a table of the two graphs: https://www.desmos.com/calculator/9uirsok0kv

OpenStudy (igreen):

For Part C, the solution when you graph it will be where the two lines intersect..

OpenStudy (kidthatbro8):

oh, alright.

OpenStudy (kidthatbro8):

for part B, what does it mean by ""

OpenStudy (igreen):

?

OpenStudy (kidthatbro8):

Take the integer values of x between -3 and 3

OpenStudy (igreen):

The table tells you..it wants you to have a table that starts with -3, and goes all the way through 3, which will be -3, -2, -1, 0, 1, 2, 3..and I already did that for you..and it shows the y-values..so you just make two tables that looks just like those two I gave you.

OpenStudy (kidthatbro8):

oh, alright !